# Tuning LLM-based Code Optimization via Meta-Prompting: An Industrial Perspective

### v1

@@ Abstract
There is a growing interest in leveraging large language models (LLMs) for automated code optimization. However, industrial platforms deploying multiple LLMs face a critical challenge: prompts optimized for one LLM often fail with others, requiring expensive model-specific prompt engineering. This cross-model prompt engineering bottleneck severely limits the practical deployment of multi-LLM optimization systems in production environments. To address this, we introduce Meta-Prompted Code Optimization (MPCO), a framework that automatically generates high-quality, task-specific prompts across diverse LLMs while maintaining industrial efficiency requirements. MPCO leverages meta-prompting to dynamically synthesize context-aware optimization prompts by integrating project metadata, task requirements, and LLM-specific contexts, and it seamlessly deploys on the Artemis industrial platform for automated validation and scaling.
@@
Our comprehensive evaluation on five real-world codebases with 366 hours of runtime benchmarking demonstrates MPCO’s effectiveness: it achieves overall performance improvements up to 19.06% with the best statistical rank across all systems compared to baseline methods. Analysis shows that 96% of the top-performing optimizations stem from meaningful edits. Through systematic ablation studies and meta-prompter sensitivity analysis, we identify that comprehensive context integration is essential for effective meta-prompting, and that all three major LLMs can serve effectively as meta-prompters, providing actionable insights for industrial practitioners.
Index Terms:
Meta-Prompting, Code Optimization, Prompt Engineering, Performance Optimization, LLM4Code, AI4SEI Introduction
@@
Large language models (LLMs) have demonstrated remarkable capabilities in many aspects across the software development lifecycle, including code completion [1], natural‑language documentation [2], test-case generation [3], automated debugging [4], and more [5]. Beyond these capabilities, recent research studies have demonstrated that LLMs can systematically analyze code patterns to uncover optimization opportunities and produce revised implementations—ranging from memory‑management enhancements and algorithmic refinements to automated parallelization and vectorization—thereby enabling high‑performance, resource‑efficient software [6, 7, 8].
@@
The success of LLM-based code optimization has led to the development of specialized industrial platforms that integrate these capabilities into production workflows. These platforms provide enterprise-grade infrastructure for deploying LLM optimization at scale, serving diverse client needs from high-frequency trading systems requiring microsecond latencies to scientific computing applications demanding maximum throughput [9, 10, 11, 12, 13]. Among others, TurinTech AI’s Artemis platform [14] represents the cutting edge of AI‑driven software performance engineering, which deploys multiple LLMs at scale to search for optimal optimizations of critical code snippets with real‑time validation, as illustrated in Figure 1. It now processes tens of thousands of requests daily, and offers automated performance‑engineering solutions that reduce development costs and accelerate time‑to‑market for performance‑critical applications [15].
@@
However, in its production deployment, Artemis revealed a fundamental obstacle in multi-LLM code optimization, which we term the “cross-model prompt engineering challenge”—a prompt that produces excellent optimizations with one LLM may yield poor results with others [16, 17]. Consequently, platforms need to develop and maintain model-specific prompts for each LLM-task combination; furthermore, industrial platforms often face strict efficiency constraints that limit the feasibility of advanced prompt optimization methods that require extensive trial-and-error or computation [17]. This raises the critical question for this work:
@@
Existing solutions cannot adequately address these requirements—conventional prompting studies like Chain-of-Thought (CoT) focus primarily on general-purpose tasks and single-model scenarios, lacking the specialized contextual awareness and cross-model adaptability required for industrial code optimization [17]; evolutionary prompt optimization methods require pre-defined evaluation functions for each project and multi-round processing, causing scalability problems [18, 19]; advanced prompt tuning agents, while more flexible and powerful, demand extensive computing resources, impractical for production environments requiring rapid processing [20, 21].
To address this, we introduce Meta-Prompted Code Optimization (MPCO), an end‑to‑end framework that automates prompt engineering while maintaining industrial requirements for flexibility and efficiency across multiple LLMs. In summary, our key contributions are:
1.
We present a novel prompt engineering framework that feeds a meta‑prompting LLM rich contextual information to efficiently generate model‑adaptive prompts, thereby addressing the cross-model prompting challenge that limits industrial platforms.
2.
@@
The remainder of this paper is structured as follows. Section II provides background and related works. Section III details the MPCO framework architecture and its integration with industrial platforms. Sections IV and V describe our comprehensive evaluation on real-world industrial codebases. We discuss the practical implications and limitations of MPCO in Section VI and conclude in Section VII.
@@ II-A Industrial Multi-LLM Code Optimization
Code optimization—the process of improving software performance via algorithmic and low‑level code transformations—has traditionally relied on handcrafted algorithmic tweaks, compiler‑driven pre‑defined transformations (e.g., loop unrolling, inlining, and dead‑code elimination), and machine learning–based autotuning systems [22]. Yet, these methods often struggle to capture domain-specific performance patterns, generalize across diverse codebases, and incur high search overheads [22, 23]. Recent advances in LLMs enable automated code optimization systems that learn from vast code corpora and leverage both static knowledge base and environment interaction to generate adaptive, context‑aware optimization suggestions, without expert intervention or limitation on pre-defined operations [6]. Today, industrial platforms integrate these LLM‑driven optimizations into production pipelines, offering low‑latency inference and high‑throughput processing [9, 14].
@@ II-B Cross-Model Prompt Engineering Bottleneck
The effectiveness of LLM‑generated optimizations depends critically on prompt quality [16, 17], yet different models exhibit varying sensitivities to prompt structure, context, and optimization strategies. As the example in Figure 2 shows, optimizing the same code snippet with three LLMs using different optimal prompt templates reveals significant performance variations. Especially, GPT-4o achieves its best performance (116.1s) with its own prompt template, but degrades by 23.5% (143.4s) when using Claude’s template and 29.0% (149.8s) with Gemini’s template. These inconsistencies stem from model‑specific factors—differences in training corpora and architectures, variations in tokenization and context‑window handling, and divergent decoding defaults and instruction‑following behaviors [24, 25, 17].
@@ II-C Existing Prompt Optimization Approaches and Limitations
Evolutionary prompting methods encompass both automatic prompt optimization approaches such as EvoPrompt (genetic algorithm-based prompt evolution) [29] and APE (automatic prompt engineering through iterative refinement) [25], as well as gradient-based prompt tuning techniques (continuous optimization of prompt embeddings) [19, 30, 16]. While these methods achieve substantial gains for well-defined tasks, they often suffer from high computational overhead and reliance on manual design of evaluation pipelines for each task/project, limiting their scalability in production environments.
Agentic prompting systems employ autonomous AI agents that iteratively refine prompts through environmental interaction, feedback incorporation, and self-reflection mechanisms [31, 32]. These agent-based approaches demonstrate context-awareness and cross-domain adaptability by dynamically adjusting prompts based on task requirements and model responses. However, they usually incur higher deployment complexity and overhead due to extensive interaction cycles, making them unsuitable for large-scale industrial scenarios requiring rapid responses.
@@ III-A Stage 1: Profiling and Bottleneck Identification
The framework begins by systematically identifying performance bottlenecks within real-world software projects. In this stage, users specify software projects requiring optimization, and the framework applies industry-standard profilers—Intel VTune Profiler [38] for C++ and Speedscope [39] for Python—to locate performance-critical code snippets. These profiling tools use instrumentation to trace function calls and resource usage at runtime, and visualize hotspots via flame graphs or detailed reports. We opt for VTune and Speedscope in this study because VTune provides low‑overhead, cycle‑accurate profiling of C++ code [40], while Speedscope offers rapid and efficient exploration for dynamic Python workloads [41].
@@ III-B Stage 2: Context Collection and Meta-Prompt Generation
To overcome the cross‑model prompt engineering bottleneck, MPCO introduces a novel meta‑prompting mechanism that swiftly collects and unifies comprehensive relevant context for prompt synthesis. Unlike prior works that either use generic templates without context awareness [37, 33] or incur large overhead in context collection [18, 32], MPCO builds on Artemis’s existing metadata store to enable rapid context gathering through three lightweight contexts:
@@ IV-C Performance Evaluation and Validation
Each generated code optimization was integrated into a new code variant, and each variant was submitted to the Artemis platform for automated compilation and execution. Any LLM response that failed to adhere to the required format (containing only the synthesized prompt/optimized code without additional commentary) was excluded from analysis. Similarly, code optimizations that resulted in compilation errors or runtime failures were systematically removed from the dataset prior to statistical analysis. This filtering process ensures that our performance comparisons are based solely on functionally correct and executable code variants, providing a fair and meaningful evaluation of optimization effectiveness across different prompting approaches.
@@
All experiments were conducted on a dedicated Google Cloud Platform server with 4 CPUs, approximately 16GB RAM, running Ubuntu 25.04. The benchmarks were executed in Docker containers to ensure consistent execution environments and eliminate interference between different optimization runs. Each experiment was executed in an isolated environment without any other processes running to minimize measurement noise and ensure reproducible results.
@@ IV-D Evaluation Metric and Statistical Analysis
To evaluate each code optimization, we adopt the most common metric in the LLM-based code optimization literature [6]—the percentage runtime performance improvement (%PI), computed as:
@@
Our ranking strategy works as follows: (1) calculate the mean %PI for each approach, (2) sort approaches by mean %PI in descending order, and (3) assign ranks sequentially—for each approach, if it is statistically significantly different from the previous approach ( or )111The significance level follows established norms in software engineering research [49], while the threshold for effect size ensures that differences considered are not only statistically significant but also practically relevant [50]., assign it the next rank; otherwise, assign it the same rank as the previous approach. By doing so, we group similar-performing methods together while distinguishing methods with meaningful performance differences.
@@ V-A1 Method
In industrial platforms, code optimizations are often expected to complete within seconds and at minimal computational cost. The advanced prompting techniques discussed in Section II, which often rely on extensive reasoning chains or agent coordination [18, 31], are thus beyond the scope of this study. Instead, we benchmarked MPCO against established baseline prompting methods most commonly used in LLM-based code optimization [6], including:
@@ V-A2 Results
Table III presents the results for RQ1. Clearly, MPCO achieves the best average rank (1.00) for all subject systems, significantly outperforming all baseline methods, and demonstrating consistent gains across the three LLMs used for code optimization.
@@
Interestingly, we can see that even when the same contexts were provided, contextual prompting did not achieve competitive results as MPCO. This highlights the strength of meta-prompting, which dynamically adjusts its guidance based on the specific task and the capabilities of the LLM. In contrast, contextual prompting may overwhelm the LLM with too much information, leading to degraded performance—similar to a teacher who uses the same teaching style for all students, regardless of their individual characteristics. In summary:
@@ V-B2 Results
Table IV highlights the importance of using comprehensive contexts to MPCO’s effectiveness, where the full MPCO template is consistently ranked the first across all systems and gets the highest %PI in three out of five, with performance gains up to 9.01%. Even in the two cases where the full MPCO template did not yield the highest %PI—Faster‑Whisper and BitmapPlusPlus—MPCO still achieved rank 1 based on statistical tests. These results demonstrate that removing any contextual component noticeably reduces MPCO’s overall effectiveness, underscoring the importance of comprehensive context integration.
For Faster‑Whisper, MPCO slightly outperformed the full template (5.73% vs. 5.64%), possibly because the optimization targets in this project are relatively well-known, making them less sensitive to model-specific instructions. In BitmapPlusPlus, all variants—including the full template and ablated versions—shared the top rank, suggesting that optimization challenges in this system are relatively straightforward, and even limited contextual guidance is sufficient, which aligns with our findings from RQ1 that performance issues in BitmapPlusPlus appeared easier to diagnose.
@@ V-C2 Results
These findings have practical implications for industrial deployment: while GPT-4o shows a slight performance advantage, all three meta-prompting LLMs provide significant improvements over the original code. Therefore, organizations can confidently choose their meta-prompting LLM based on practical considerations—such as cost, availability, and platform compatibility—without significantly compromising MPCO’s effectiveness. Consequently, we conclude:
@@ VI-A Practical Validity of MPCO Performance Improvements
Our comprehensive evaluation across five real-world systems demonstrates that MPCO’s performance improvements are both significant and robust. However, a critical concern in automated code optimization is whether performance improvements represent genuine algorithmic enhancements rather than just removing the original functions.
@@
Notably, Loop & Vectorization (26%), Algorithmic Improvements (20%), and Code Quality & Correctness improvements (20%) are the most targeted optimizations, whereas the small proportion (4%) of trivial optimizations—such as reducing the number of iterations and adjusting line breaks—indicates that MPCO effectively identifies common performance bottlenecks [27, 6] and applies appropriate optimization strategies.
VI-B Actionable Guidelines for Context Provision

### v2

@@ Abstract
There is a growing interest in leveraging multiple large language models (LLMs) for automated code optimization. However, industrial platforms deploying multiple LLMs face a critical challenge: prompts optimized for one LLM often fail with others, requiring expensive model-specific prompt engineering. This cross-model prompt engineering bottleneck severely limits the practical deployment of multi-LLM systems in production environments. We introduce Meta-Prompted Code Optimization (Mpco), a framework that automatically generates high-quality, task-specific prompts across diverse LLMs while maintaining industrial efficiency requirements. Mpco leverages meta-prompting to dynamically synthesize context-aware optimization prompts by integrating project metadata, task requirements, and LLM-specific contexts. It is an essential part of the Artemis code optimization platform for automated validation and scaling.
@@
Our comprehensive evaluation on five real-world codebases with 366 hours of runtime benchmarking demonstrates Mpco’s effectiveness: it achieves overall performance improvements up to 19.06% with the best statistical rank across all systems compared to baseline methods. Analysis shows that 96% of the top-performing optimizations stem from meaningful edits. Through systematic ablation studies and meta-prompter sensitivity analysis, we identify that comprehensive context integration is essential for effective meta-prompting and that major LLMs can serve effectively as meta-prompters, providing actionable insights for industrial practitioners.
I Introduction
@@
Large language models (LLMs) have demonstrated remarkable capabilities in many aspects across the software development lifecycle, including code completion [1], natural‑language documentation [2], test-case generation [3], automated debugging [4], and more [5]. Beyond these capabilities, recent research studies have demonstrated that LLMs can systematically analyze code patterns to uncover optimization opportunities and produce revised implementations, ranging from memory-management enhancements and algorithmic refinements to automated parallelization and vectorization—thereby enabling high‑performance, resource‑efficient software [6, 7, 8].
@@
The success of LLM-based code optimization has led to the development of industrial solutions that integrate these capabilities into production workflows. These platforms provide enterprise-grade infrastructure for deploying LLM-driven code optimization at scale, serving diverse client needs from high-frequency trading systems requiring microsecond latencies to scientific computing applications demanding maximum throughput [9, 10, 11, 12, 13]. Among others, TurinTech AI’s Artemis platform [14] represents a leading AI-driven software performance engineering system, which deploys multiple LLMs at scale to search for optimal optimizations of critical code snippets with real-time validation, as illustrated in Figure 1. It now processes tens of thousands of requests daily and offers automated performance-engineering solutions that reduce development costs and accelerate time-to-market for performance-critical applications [15].
@@
However, in its production deployment, Artemis exposes a fundamental obstacle in multi-LLM code optimization, which we term the “cross-model prompt engineering challenge” - a prompt that produces excellent optimizations with one LLM may yield poor results with others [16, 17]. Consequently, platforms need to develop and maintain model-specific prompts for each LLM-task combination; furthermore, industrial platforms often face strict efficiency constraints that limit the feasibility of advanced prompt optimization methods that require extensive trial-and-error or computation [17]. This raises the critical question for this work:
@@
Unfortunately, existing solutions cannot adequately address these requirements. Conventional prompting studies like Chain-of-Thought (CoT) focus primarily on general-purpose tasks and single-model scenarios, lacking the specialized contextual awareness and cross-model adaptability required for industrial code optimization [17]. Evolutionary prompt optimization methods require pre-defined evaluation functions for each project and multi-round processing, causing scalability problems [18, 19]. Advanced prompt tuning agents, while more flexible and powerful, demand extensive computing resources, impractical for production environments requiring rapid processing [20, 21].
To address this, we introduce Meta-Prompted Code Optimization (Mpco), an end-to-end framework that automates prompt engineering while maintaining industrial requirements for flexibility and efficiency across multiple LLMs. In summary, our key contributions are:
We present a novel prompt engineering framework that feeds a meta-prompting LLM rich contextual information to efficiently generate model-adaptive prompts, thereby addressing the cross-model prompting challenge that limits industrial platforms.
@@ II-A Industrial Multi-LLM Code Optimization
Code optimization aims to improve software performance (e.g., making program run fast or reducing the binary size) via algorithmic and low-level code transformations. It has traditionally relied on handcrafted algorithmic tweaks, compiler-driven pre-defined transformations (e.g., loop unrolling, inlining, and dead-code elimination), and machine learning–based autotuning systems [22]. Yet, these methods often struggle to capture domain-specific performance patterns, generalize across diverse codebases, and incur high search overheads [22, 23]. Recent advances in LLMs enable automated code optimization systems that learn from vast code corpora and leverage both a static knowledge base and environment interaction to generate adaptive, context‑aware optimization suggestions, without expert intervention or limitation on pre-defined operations [6]. Today, industrial platforms integrate these LLM-driven optimizations into production pipelines, offering low‑latency inference and high-throughput processing [9, 14].
@@ II-B Cross-Model Prompt Engineering Bottleneck
The effectiveness of LLM‑generated optimizations depends critically on prompt quality [16, 17], yet different models exhibit varying sensitivities to prompt structure, context, and optimization strategies. As a motivating example, consider optimizing a hot code region of the Llama.cpp inference framework [24]. Figure 2 shows that optimizing the same code snippet with three LLMs using different optimal prompt templates reveals significant performance variations. Especially, GPT-4o achieves its best performance (116.1s) with its own prompt template, but degrades by 23.5% (143.4s) when using Claude’s template and 29.0% (149.8s) with Gemini’s template. These inconsistencies often stem from model-specific factors - differences in training corpora and architectures, variations in tokenization and context-window handling, and divergent decoding defaults and instruction‑following behaviors [25, 26, 17].
@@
This challenge is uniquely severe in code optimization compared to generic natural language processing (NLP) tasks. Unlike text generation, where outputs can be stylistically varied, optimized code has strict, non-negotiable correctness requirements: it must compile, pass all functional tests, and be free of runtime side effects like memory leaks. The determinism and reproducibility of optimizations are paramount. A slight, seemingly innocuous change in a prompt can lead to a syntactically correct but functionally flawed or slower implementation, making the cross-model challenge a critical barrier to reliable industrial deployment.
@@ II-C Existing Prompt Optimization Approaches and Limitations
Evolutionary prompting methods encompass both automatic prompt optimization approaches such as EvoPrompt (genetic algorithm-based prompt evolution) [18] and APE (automatic prompt engineering through iterative refinement) [26], as well as gradient-based prompt tuning techniques (continuous optimization of prompt embeddings) [19, 30, 16]. While these methods achieve substantial gains for well-defined tasks, they require multiple rounds of LLM interaction, leading to high computation overhead [18], which could be infeasible for platforms processing tens of thousands of daily requests [15]. Furthermore, they often rely on manual design of evaluation pipelines for each task/project, limiting their scalability in production environments.
Agentic prompting systems employ autonomous AI agents that iteratively refine prompts through environmental interaction, feedback incorporation, and self-reflection mechanisms [31, 32]. These agent-based approaches demonstrate context-awareness and cross-domain adaptability by dynamically adjusting prompts based on task requirements and model responses. However, they usually incur higher deployment complexity and overhead due to extensive environment interaction cycles, action planning, agent state updates, and memory I/O [33], making them unsuitable for large-scale industrial scenarios requiring rapid responses.
@@ III-A Stage 1: Profiling and Bottleneck Identification
The framework begins by systematically identifying performance bottlenecks within real-world software projects. In this stage, users specify software projects requiring optimization, and the framework applies industry-standard profilers - Intel VTune Profiler [39] for C++ and Speedscope [40] for Python - to locate performance-critical code snippets. These profiling tools use instrumentation to trace function calls and resource usage at runtime, and visualize hotspots via flame graphs or detailed reports. We opt for VTune and Speedscope in this study because VTune provides low‑overhead, cycle‑accurate profiling of C++ code [41], while Speedscope offers rapid and efficient exploration for dynamic Python workloads [42].
@@ III-B Stage 2: Context Collection and Meta-Prompt Generation
To overcome the cross‑model prompt engineering bottleneck, Mpco introduces a novel meta‑prompting mechanism that swiftly collects and unifies comprehensive relevant context for prompt synthesis. Unlike prior works that either use generic templates without context awareness [38, 34] or incur large overhead in context collection [18, 32], Mpco builds on Artemis’ existing metadata store to enable rapid context gathering through three lightweight contexts:
@@ IV-C Performance Evaluation and Validation
Each generated code optimization was integrated into a new code variant, and each variant was submitted to Artemis for automated compilation and execution. Any LLM response that failed to adhere to the required format (containing only the synthesized prompt/optimized code without additional commentary) was excluded from analysis. Similarly, code optimizations that resulted in compilation errors or runtime failures were systematically removed from the dataset prior to statistical analysis. This filtering process ensures that our performance comparisons are based solely on functionally correct and executable code variants, providing a fair and meaningful evaluation of optimization effectiveness across different prompting approaches.
@@
All experiments were conducted on a dedicated Google Cloud Platform server with 4 CPUs, 16GB RAM, running Ubuntu 25.04. The benchmarks were executed in Docker containers to ensure consistent execution environments and eliminate interference between different optimization runs. Each experiment was executed in an isolated environment without any other processes running to minimize measurement noise and ensure reproducible results.
@@ IV-D Evaluation Metric and Statistical Analysis
To evaluate each code optimization, we adopt the most common metric in the LLM-based code optimization literature [6], i.e., the percentage runtime performance improvement (%PI), computed as:
@@
Our ranking strategy works as follows: (1) calculate the mean %PI for each approach, (2) sort approaches by mean %PI in descending order, and (3) assign ranks sequentially—for each approach, if it is statistically significantly different from the previous approach ( or )111The significance level follows established norms in SE research [50], while the threshold for effect size ensures that differences considered are not only statistically significant but also practically relevant [51]., assign it the next rank; otherwise, assign it the same rank as the previous approach. By doing so, we group similar-performing methods together while distinguishing methods with meaningful performance differences.
@@ V-A1 Method
In industrial platforms, code optimizations are often expected to complete within seconds and at minimal computational cost. The advanced prompting techniques discussed in Section II, which often rely on extensive reasoning chains or agent coordination [18, 31], are thus beyond the scope of this study. Instead, we benchmarked Mpco against established baseline prompting methods most commonly used in LLM-based code optimization [6]222The detailed implementation of all baseline prompts is available here., including:
@@ V-A2 Results
Table III presents the results for RQ1. Clearly, Mpco achieves the best average rank (1.00) for all subject systems, significantly outperforming all baseline methods and demonstrating consistent gains across the three LLMs used for code optimization.
@@
Interestingly, we can see that even when the same contexts were provided, contextual prompting did not achieve competitive results as Mpco. This highlights the strength of meta-prompting, which dynamically adjusts its guidance based on the specific task and the capabilities of the LLM. In contrast, contextual prompting may overwhelm the LLM with too much information, leading to degraded performance - similar to a teacher who uses the same teaching style for all students, regardless of their individual characteristics. In summary:
@@ V-B2 Results
Table IV highlights the importance of using comprehensive contexts to Mpco’s effectiveness, where the full Mpco template is consistently ranked the first across all systems and gets the highest %PI in three out of five, with performance gains up to 9.01%. Even in the two cases where the full Mpco template did not yield the highest %PI - Faster‑Whisper and BitmapPlusPlus - Mpco still achieved rank 1 based on statistical tests. These results demonstrate that removing any contextual component noticeably reduces Mpco’s overall effectiveness, underscoring the importance of comprehensive context integration.
For Faster‑Whisper, Mpco slightly outperformed the full template (5.73% vs. 5.64%), possibly because the optimization targets in this project are relatively well-known, making them less sensitive to model-specific instructions. In BitmapPlusPlus, all variants, including the full template and ablated versions, shared the top rank, suggesting that optimization challenges in this system are relatively straightforward, and even limited contextual guidance is sufficient, which aligns with our findings from RQ1 that performance issues in BitmapPlusPlus appeared easier to diagnose.
@@ V-C2 Results
These findings have practical implications for industrial deployment: while GPT-4o shows a slight performance advantage, all three meta-prompting LLMs provide significant improvements over the original code. Therefore, organizations can confidently choose their meta-prompting LLM based on practical considerations, such as cost, availability, and platform compatibility, without significantly compromising Mpco’s effectiveness. Consequently, we conclude:
@@ VI-A Practical Validity of Mpco Performance Improvements
Our evaluation across five real-world systems demonstrates that Mpco’s performance improvements are both significant and robust. However, a critical concern in automated code optimization is whether performance improvements represent genuine algorithmic enhancements rather than just removing the original functions.
@@
Notably, Loop & Vectorization (26%), Algorithmic Improvements (20%), and Code Quality & Correctness improvements (20%) are the most targeted optimizations, whereas the small proportion (4%) of trivial optimizations, such as reducing the number of iterations and adjusting line breaks, indicates that Mpco effectively identifies common performance bottlenecks [28, 6] and applies appropriate optimization strategies.
VI-B Illustrative Examples of Generated Meta-Prompts
To illustrate Mpco’s superiority over baseline methods, we examine concrete examples from Llama.cpp, where Mpco achieved 97.5% higher %PI over the best baseline (7.87 vs 3.97 in Table III).
Particularly, Mpco generates prompts with comprehensive optimization strategies, while adapting to each LLM’s strengths (Appendix VIII). For example: GPT-4o (Figure 7) focuses on “complex interdependencies and comprehensive optimization”; Claude 3.7 Sonnet (Figure 8) promotes “systematic architectural thinking with maintainability considerations; and Gemini 2.5 Pro (Figure 9) emphasizes “complex reasoning to verify performance metrics”.
In contrast, baseline methods provide generic guidance: chain-of-thought offers step-by-step frameworks without domain expertise (Figure 10); few-shot learning provides simple examples insufficient for C/C++ optimization (Figure 11); and contextual prompting uses all available contexts (project, task, LLM) uniformly across all cases (Figure 12).
Therefore, the key advantage of Mpco lies in its dual adaptation strategy: (1) task-specific guidance that addresses the unique characteristics of Llama.cpp (C/C++ codebase, performance-critical AI inference, memory-intensive operations), and (2) LLM-specific instructions tailored to each model’s cognitive strengths and architectural capabilities. This explains why Mpco achieved top-ranked performance across all systems in RQ1, while baseline methods failed to consistently leverage each LLM’s optimization capabilities or address task-specific performance bottlenecks.
VI-C Meta-Prompting Overhead Analysis
To validate Mpco’s industrial efficiency claims, we conducted a comprehensive overhead analysis by measuring meta-prompting overhead across 50 constructs from all five subject systems, using GPT-4o as the meta-prompter. Figure 6 presents the detailed timing breakdown and token usage analysis.
The results demonstrate that Mpco introduces minimal computational overhead: averaging 3.8 seconds per meta-prompt generation with 624 tokens (428 input, 196 output). In particular, the LLM query dominates the timing, while context collection, template filling, and response processing contribute negligible overhead (0.1ms each). At GPT-4o pricing ($2.5/1M input tokens, $10/1M output tokens), each meta-prompt generation costs approximately $0.003, making the approach highly practical for industrial deployment.
Importantly, Mpco provides substantial cost advantages over traditional prompt engineering approaches. Manual prompt crafting requires expert engineers to spend minutes to hours developing and tuning LLM-specific prompts for each optimization task, followed by ongoing maintenance as LLMs evolve. In contrast, Mpco automates this entire process at $0.003 per LLM-task pair, orders of magnitude lower cost than evolutionary prompt optimization or iterative agent-based methods that require multiple LLM calls. The single-call design ensures that once generated, meta-prompts can be reused indefinitely without additional overhead, reducing both the expertise requirement and maintenance burden while delivering superior optimization effectiveness across diverse tasks and LLMs.
VI-D Actionable Guidelines for Context Provision